Recently, Experian reported that the average credit score for a new-car loan was 753. Suppose Ally Financial, a bank holding company that finances car loans, would like to test the hypothesis that the average credit score has increased since the Experian report. A random sample of 20 new-car loans had an average credit score of 764.2 with a sample standard deviation of 34.5. Ally Financial would like to set α = 0.05. The critical value for this hypothesis test would be ________.

a. 1.282

b. 2.977

c. 1.729

d. 2.080
Answer #1

solution:

Degrees of freedom = df = n - 1 = 20 - 1 = 19

=  0.05

t ,df = t0.05,19 =1.729 ( using student t table)